OpenStudy (anonymous):

FAN AND MEDAL: The equation below represents Function A and the graph represents Function B: Function A f(x) = x - 9 Function B graph of line going through ordered pairs negative 1, negative 3 and 2, 3 Which equation best compares the slopes of the two functions? Slope of Function B = 2 × Slope of Function A. Slope of Function A = Slope of Function B Slope of Function A = 2 × Slope of Function B Slope of Function B = - Slope of Function A

OpenStudy (anonymous):

@Jaydenthequiet

OpenStudy (shamim):

y=x-9

OpenStudy (jaydenthequiet):

function B equals 1/2

OpenStudy (shamim):

can u tell me the slope of this line

OpenStudy (anonymous):

@shamim Which one is that?

OpenStudy (anonymous):

idk...i'm not sure how to do any of this...

OpenStudy (shamim):

let f(x)=y

OpenStudy (shamim):

the standard equation of a line is y=mx+c

OpenStudy (shamim):

m=?

OpenStudy (shamim):

use x=x*1

OpenStudy (anonymous):

Doesn't M = the slope? I think..?

OpenStudy (shamim):

slope=m

OpenStudy (anonymous):

Okay, so how does that help me here..?

OpenStudy (anonymous):

Where do I go from here?

OpenStudy (shamim):

first i m trying to find out 2 slopes frm 2 rquation of line.

OpenStudy (shamim):

then i will find out the relationship between those two slopes

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

Well, how do I do that?

OpenStudy (shamim):

anywsy y=mx+c y=x-9 just compare the above two equations

OpenStudy (anonymous):

@Courtneydyrdek98

OpenStudy (shamim):

m=1?

OpenStudy (anonymous):

The second one has no m..?

OpenStudy (anonymous):

So is it A?

OpenStudy (shamim):

we will try to find out m frm second equation

OpenStudy (anonymous):

2?

OpenStudy (shamim):

why u r just trying to get the result?

OpenStudy (anonymous):

I don't understand how to do any of this...Like, it's REALLY confusing. I don't know how to do slope or any of this...

OpenStudy (shamim):

ok

OpenStudy (anonymous):

Can you please explain it in a different way?

OpenStudy (shamim):

do u know a equation of a line is y=mx+c

OpenStudy (anonymous):

I thought it was y=mx+b ?

OpenStudy (shamim):

response plz

OpenStudy (shamim):

its not really difficuly

OpenStudy (shamim):

easy to understand

OpenStudy (anonymous):

What is your question?

OpenStudy (anonymous):

it isn't easy for me...that's why I'm asking for your help..